A rocket blasts off from rest and attains a speed of 45 m/s in 15 sec. An astronaut has a mass of 57 kg. What is the astronaut's apparent weight during takeoff?
1
Expert's answer
2020-03-06T10:24:45-0500

As per the question,

Initial velocity of the rocket is (u)=0

Final velocity (v)=45 m/sec

time (t)=15 sec

Mass of the astronaut =57 Kg

Now, applying the newton's first law of motion,

Let the acceleration of the rocket is a,

"v=u+at"

"\Rightarrow 45=0+a\times 15"

"\Rightarrow a=\dfrac{45}{15}=3 m/sec^2"

Now,

During the take off of the rocket, pseudo force will work on astronaut.

Hence, "F=ma =57\times 3 =171N"

Hence total reading of the weighting machine ="mg+ma"

"=171+57\times 9.8 =729.6N"

Hence total reading of the weighting machine will be 729.6N.